diff eq
the problem states that to solve the given linear initial-value problem use the power series method.
please include intermediate steps
(x² - 1)y+ 3xy + xy = 0, y(0) =4 , y(0) = 6
Answer #1

, y lol = 4 - 7 10) = 6 Q solve Jup by hower series method [22-1) y + 3x y + xy =0 - 0 som Tom comparing std. diff egn P/Xgiven that Y(0) = 4 .. a lot a,(0) + a20 = yo) 4 = a + => 4 = ao y lol = 6 2 9(%) : 8 - | a, + Q. 3, (0) + 3, 0° 6 = a, the= (m - 1) - 2 => S - 3 n + 1)(x + 2) a + 2 1 nel na + 230 am 2 + * = 0 => (m) am X -[204 +6052 + en av mine smrti) +[30,2& mem-v) am 1 nln-il an ante = -lna)/n+2) On+ 2 + 3 anant any = 0 1 l mln-il an + 3 nam tam., 1 4) (+2) > chillen we in pnu 2